OpenStudy (mysticfury):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y = x + Line 2: y = x +

Mehek (mehek14):

do you know the slope formula?

OpenStudy (mysticfury):

or a screenshot

OpenStudy (mysticfury):

yes y = mx + b

Mehek (mehek14):

no the slope formula like how to find `m`?

OpenStudy (mysticfury):

oh yeah x1 - y1 x2 - y2

OpenStudy (mysticfury):

wait i did that wrong lol >.<

Mehek (mehek14):

nope \(\bf\dfrac{y_1-y_2}{x_1-x_2}\)

OpenStudy (mysticfury):

thats what i meant ^^

Mehek (mehek14):

or \(\bf\dfrac{y_2-y_1}{x_2-x_1}\)

Mehek (mehek14):

both of those work

OpenStudy (mysticfury):

ok so ill do that, and you can check my answer for the slopes, okay?

Mehek (mehek14):

sure

OpenStudy (mysticfury):

okay i got 2/1 for the first one 2/-1 for the second one

OpenStudy (mysticfury):

the y intercepts are on top and the x intercepts are on the bottom ^^^^

OpenStudy (mysticfury):

@Qwertty123 @Mehek14

Mehek (mehek14):

wdym ?

OpenStudy (mysticfury):

?

Mehek (mehek14):

the slope for both lines is 2

OpenStudy (mysticfury):

alrighty then how do i find the y intercept for both lines?

Mehek (mehek14):

first one already tells you when x = 0 y= ?

OpenStudy (mysticfury):

3? but what about the second set of numbers

OpenStudy (mysticfury):

under neath it i ment

OpenStudy (mysticfury):

1 and 5

OpenStudy (mysticfury):

and about the slope this "Use a decimal for the slope if necessary." makes me think it should be a decimal :/

Mehek (mehek14):

slope is 3 for both

Mehek (mehek14):

y-intercept***

Mehek (mehek14):

i'm honestly lost why are both the same

OpenStudy (mysticfury):

okay so my answer for both is y = 2x + 3 ??

OpenStudy (mysticfury):

idk it dosent really make sense

Mehek (mehek14):

second line is y = 2x + 3

Mehek (mehek14):

yep I checked both are the same lines

OpenStudy (mysticfury):

okay ill try it then

OpenStudy (mysticfury):

it was correct.

Mehek (mehek14):

ok good :)

OpenStudy (mysticfury):

thanks ...

Mehek (mehek14):

yw

OpenStudy (qwertty123):

Nice Job :)